x= adult tickets y= children tickets
5x + 3y = 134
x + y = 32
whats true about this equation?
a.) More adults attended the reunion than children
b.) six more children attended that adults
c.) nineteen children attended the reunion
d.) three more adults attended than children
i didnt know how i would solve this?

Answer by mananth(16946)

5x + 3y = 134
x + y = 32. multiply by -3
-3x-3y=-96
add to the first equation
we get

2x=38
x=19
if x=17 plug x in x+y=32
y=13

Adults = 19
children = 13
now you may find the correct statement